OpenStudy (lilmike234):

What expression is equivalent to 4x^3+2x^2+x+3/x+2 ?

OpenStudy (alekos):

whats the rest of the question? we're missing something

OpenStudy (lilmike234):

That's the whole question. The multiple choice consist of 1. 4x^2-6x+13-23/x+2 2. 4x-6+13/x+2 3. 4x^3-6x+13+23/x+2 4.4x^2+13-2/x+2

OpenStudy (anonymous):

\[\frac{4 x^3+2 x^2+x+3}{x+2}=4 x^2-6 x+13-\frac{23}{x+2} \]

OpenStudy (alekos):

I thought there was something missing! thanks @robtobey